Puan Siti Lela is a Malaysian and plans to visit Paris or New York to buy a Louis
Vuitton (LV) bag that cost uro $30,000 in Paris and USD28,000 in New York.
Assuming no cost in transfers or transportation. Which is a cheaper destination for
Puan Siti Lela to buy her LV handbag by showing your conversion calculation? How
many uros or USD does Puan Siti Lela need to pay for the LV handbag?
1
Expert's answer
2020-07-08T17:48:27-0400

Current exchange rate is:

1 USD = 0.89 euro or 1 euro = 1.13 USD.

So, €30,000 = 1.13×30,000 = $33,900 or $28,000 = 0.89×28,000 = €24,920.

That's why Puan Siti Lela should go to New York, where LV handbag is cheaper.